Suppose that pulse rates among healthy adults are normally distributed with a mean of 79 beats/minute and a standard deviation of 8 beats/minute. What proportion of healthy adults have pulse rates that are more than 81 beats/minute? Round your answer to at least four decimal places.
Answer #1

µ = 79

sd = 8

X_μ 81

                         =P(Z>\frac{81-79}{8})

                         = P(Z > 0.25)

                         = 1 - P(Z < 0.25)

                         = 1 - 0.5987

                         = 0.4013 (ans)